# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v4

@@ A Proofs
where $H_{\alpha\alpha^{\prime}}\left(\theta\right)$ is $\left(N-1\right)K\times\left(N-1\right)K$, $H_{\gamma\gamma^{\prime}}\left(\theta\right)$ is $\left(T-1\right)K\times\left(T-1\right)K$, $H_{\alpha\gamma^{\prime}}\left(\theta\right)$ is $\left(N-1\right)K\times\left(T-1\right)K$, and $H_{\gamma\alpha^{\prime}}\left(\theta\right)=\left[H_{\alpha\gamma^{\prime}}\left(\theta\right)\right]^{\prime}$. The next lemma states that the quadratic form $\mathbb{E}\left\{s^{\prime}(\theta)[{H}\left(\theta\right)]^{-1}s(\theta)\right\}$ can be decomposed into two dominating terms of order $O(1/\sqrt{NT})$. They establish the two additive bias terms $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$, arising from estimating individual effects and time effects, respectively.
Lemma A.1
@@
$=-\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)+2\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right](1+o_{\mathbb{P}}(1)).$ (A.1)
@@
$\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right]=O\left(1/(NT)\right),$ (A.2)
@@
$\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right]=O(1/T),\qquad\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right]=O(1/N).$ (A.3)
@@
First, under Assumptions 1 and 1, and by Davydov’s inequality, it follows that $\sup_{i,j}\sup_{\theta\in\Theta}\operatorname*{Var}[h_{i,j}^{\alpha}(\theta)]=o(1)$ as $N,T\rightarrow\infty$, where $\operatorname*{Var}$ denotes the variance and $h_{i,j}^{\alpha}(\theta)$ is the $\left(i,j\right)$-th element of $N\cdot H_{\alpha\alpha^{\prime}}\left(\theta\right)$ with $i,j=1,\ldots,N-1$. Invoking the Markov’s inequality,
@@
uniformly for $\theta\in\Theta$, where $\|A\|_{\max}$ denotes the max norm of a matrix $A$. In addition, by matrix inversion in block form, it follows that
@@
For $i=1,\ldots,N-1$ and $t=1,\ldots,T-1$, define $h_{i,t}^{\ast\alpha\gamma}(\theta)$ as the $(i,t)$-th element of $H_{\alpha\gamma^{\prime}}^{\ast}(\theta)$. We first show (A.2). We have
@@
$=:A_{NT,1}(\theta)-A_{NT,2}(\theta)-A_{NT,3}(\theta)+A_{NT,4}(\theta).$ (A.4)
@@
×$\quad\quad\times\sup\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\left[\left(1-\frac{m}{T-1}\right)[a_{i}(m)]^{1-2/\eta}\right]+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
@@
×$\quad\quad\times\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\left[\left(1-\frac{m}{T-1}\right)[a_{i}(m)]^{1-2/\eta}\right]+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
@@
Therefore, (A.2) follows from (A.4).
@@
noting that $\sup_{i\neq j}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in $\theta$. Thus, the first part of (A.3) follows from (A.5). The second part of (A.3) follows from a similar procedure.
Combining (A.1), (A.2), and (A.3),
@@
$r^{s}\left(\theta,a\right)$ $:=\left(r_{\psi_{1}}^{s}\left(\theta,a_{1}\right),\ldots,r_{\psi_{(N+T-2)K}}^{s}(\theta,a_{(N+T-2)K})\right)^{\prime},$ (A.6)
@@
Proof. In the proof, we retain the notation and setup from Lemma A.1. From (2.13), it is evident that the dominant term in the expansion of $\widehat{\psi}(\theta)-\psi\left(\theta\right)$ is $[-{H}(\theta)]^{-1}s(\theta)$. Specifically,
@@
$=-\begin{pmatrix}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\\ H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\end{pmatrix}+[-\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\cdot o_{\mathbb{P}}(1)$ (A.7)
@@
$N\sqrt{T}u_{i}^{\alpha}(\theta)=N\sum_{j=1}^{N-1}\left[h_{i,j}^{\ast\alpha}(\theta)\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{j,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)\right].$ (A.8)
@@
$T\sqrt{N}u_{i}^{\alpha\gamma}(\theta)=\sum_{t=1}^{T-1}\left[h_{i,t}^{\ast\alpha\gamma}(\theta)\left(\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\left[\widetilde{s}_{j,t}^{\gamma}(\theta)-\widetilde{s}_{j,T}^{\gamma}(\theta)\right]\right)\right].$ (A.9)
@@
uniformly in $\theta$. From (A.8) and (A.9), we conclude that
$\sup_{\theta\in\Theta}\|H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(N\sqrt{T})\right),\qquad\sup_{\theta\in\Theta}\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(T\sqrt{N})\right).$ (A.10)
@@
$\sup_{\theta\in\Theta}\|H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(T\sqrt{N})\right),\qquad\sup_{\theta\in\Theta}\|H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(N\sqrt{T})\right).$ (A.11)
Thus, combining (A.7), (A.10), and (A.11), we have
@@
$\left\|\widehat{\psi}(\theta)-\psi(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\left(1/(NT)^{3/4}\right).$ (A.12)
From (A.12), we conclude that $\sup_{\theta\in\Theta}\|G_{\psi_{j}}\left(\theta,a_{j}\right)-\partial_{\psi_{j}}H(\theta)\|_{\max}=o_{\mathbb{P}}(1)$. Without loss of generality, let $\psi_{j}=\alpha_{i}$ for $i=1,\ldots,N-1$. We can express $\partial_{\alpha_{i}}H(\theta)$ as
@@
where $e_{i}$ is an $(N-1)\times 1$ vector with the $i$-th component equal to $1$ and all other elements equal to 00. Using (A.12), for each $j=1,\ldots,N-1$, we write
@@
$=:\frac{1}{2}\left\{C_{NT,1}(\theta)+C_{NT,2}(\theta)+C_{NT,3}(\theta)\right\}\left[1+o_{\mathbb{P}}(1)\right].$ (A.13)
Let $v_{i}^{\alpha}(\theta)$ denote the $i$-th element of $H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)$, for $i=1,\ldots,N-1$. Using (A.8) and (A.9), it can be shown that
@@
$=O_{\mathbb{P}}\left(1/(NT)\right).$ (A.14)
@@
$=O_{\mathbb{P}}\left(1/(NT)\right),$ (A.15)
@@
$r(\theta)$ $:=\frac{1}{2}(r^{s}(\theta,a))^{\prime}[H\left(\theta\right)]^{-1}r^{s}(\theta,a)+r^{l}(\theta,a_{0}),$ (A.16)
$r^{l}\left(\theta,a_{0}\right)$ $:=\frac{1}{3}\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}(\theta,a_{0}\times\iota_{(N+T-2)K})$ (A.17)
@@
Lemma A.3
Under Assumption 1, $r(\theta)$ defined by (A.16) satisfies
@@
Proof. In the proof, we retain the notation and setup from Lemmas A.1 and A.2. From Lemma A.2, we know that $NT\cdot r^{s}\left(\theta,a\right)$ is comparable to $(NT)^{3/4}\cdot s(\theta)$. Following similar arguments as in Lemma A.1, we can show that
@@
Moreover, from (A.12), (A.14), and (A.14), we have
@@
$=O_{\mathbb{P}}\left(1/(NT)^{5/4}\right).$ (A.18)
By Assumption 1 and applying the $L_{1}$ convergence theorem to A.18, we find that
@@
for $r\left(\theta\right)$ defined in (A.16). The claim in Theorem 1 follows directly from Lemmas A.1 and A.3.
@@
Proof of Theorem 2. We retain the notation introduced in Lemmas A.1-A.3. We first show
$\widehat{B}_{\alpha}(\theta)-B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),\qquad\widehat{B}_{\gamma}(\theta)-B_{\gamma}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$ (A.19)
To show the first part of (A.19), note that
@@
$S_{\alpha\alpha}^{1}\left(\theta\right)$ $:=\operatorname*{diag}\left\{\widehat{s}_{1}^{\alpha}\left(\theta\right)\left(\widehat{s}_{1}^{\alpha}\left(\theta\right)\right)^{\prime},\ldots,\widehat{s}_{N-1}^{\alpha}\left(\theta\right)\left(\widehat{s}_{N-1}^{\alpha}\left(\theta\right)\right)^{\prime}\right\},$
(${}_{\alpha\alpha}^{2}\left(\theta\right)$ $:=(\iota_{N-1}\iota_{N-1}^{\prime})\otimes\left[\widehat{s}_{N}^{\alpha}\left(\theta\right)\left(\widehat{s}_{N}^{\alpha}\left(\theta\right)\right)^{\prime}\right].$
@@
×$\quad\quad\times\operatorname*{diag}\left\{\widetilde{s}_{1,t}^{\alpha}(\theta)\left(\widetilde{s}_{1,s}^{\alpha}(\theta)\right)^{\prime}-\widehat{s}_{1,t}^{\alpha}(\theta)\left(\widehat{s}_{1,s}^{\alpha}(\theta)\right)^{\prime},\ldots,\widetilde{s}_{N-1,t}^{\alpha}(\theta)\left(\widetilde{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}-\widehat{s}_{N-1,t}^{\alpha}(\theta)\left(\widehat{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}\right\}$
@@
×$\quad\quad\times\operatorname*{diag}\left\{\mathbb{E}\left[\widetilde{s}_{1,t}^{\alpha}(\theta)\left(\widetilde{s}_{1,s}^{\alpha}(\theta)\right)^{\prime}\right],\ldots,\mathbb{E}\left[\widetilde{s}_{N-1,t}^{\alpha}(\theta)\left(\widetilde{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}\right]\right\}.$ (A.20)
@@
given that $\tau/T\rightarrow 0$. Combining this with (A.20), we have
$\sup_{\theta\in\Theta}N^{2}\left\|S_{\alpha\alpha}^{1}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)\right\|_{\max}=o_{\mathbb{P}}(1/T).$ (A.21)
From the proof of Lemma A.2, we also have $\sup_{\theta\in\Theta}\|\widehat{\psi}(\theta)-\psi(\theta)\|_{\max}=o_{\mathbb{P}}(1)$. Moreover, under the conditional strong mixing assumption, each element of $T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ converges in probability to the corresponding element in $T^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ by the law of large numbers. Therefore (A.21) further leads to
@@
where the last equality follows from, as shown in Lemma A.2, $\sum_{j=1}^{N-1}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta$. The above leads to
$\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)\left[S_{\alpha\alpha}\left(\theta\right)-\widehat{S}_{\alpha\alpha}\left(\theta\right)\right]\right\}=o_{\mathbb{P}}(1/\sqrt{NT}).$ (A.22)
@@
$\operatorname*{trace}\left\{\left[\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)-H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\right]S_{\alpha\alpha}\left(\theta\right)\right\}=o_{\mathbb{P}}(1)\cdot\operatorname*{trace}\left\{H_{\alpha\alpha^{\prime}}^{\ast}(\theta)S_{\alpha\alpha}\left(\theta\right)\right\}=o_{\mathbb{P}}(1)\cdot B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$ (A.23)
Combining A.22 and A.23 leads to the first part of A.19. The second part of A.19 can be shown similarly.
@@
As from Lemma A.1, we have
@@
The conclusion follows by combining this and (A.19).
@@
$0=\triangledown_{\theta}l(\theta_{0})+\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0})+o_{\mathbb{P}}(1/\sqrt{NT})$ (A.24)
@@
Thus, using Assumption 2 and Slutsky’s theorem, we obtain from (A.24) that
$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})=-\left\{\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\right\}^{-1}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.25)
Given the conditional cross-sectional independence and the conditional strong mixing condition (Assumption 1), (A.25) leads to
$\left[\Sigma^{-1}\Omega\Sigma^{-1}\right]^{-1/2}\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}\left(0,\mathbb{I}_{K}\right)$ (A.26)
@@
Since we consider a likelihood model, $\Sigma^{-1}\Omega\Sigma^{-1}$ in (A.26) simplifies to $\Sigma^{-1}$ under the information matrix equality ($\Sigma=\Omega$).

### v5

@@ E Proofs
where $H_{\alpha\alpha^{\prime}}\left(\theta\right)$ is $\left(N-1\right)K\times\left(N-1\right)K$, $H_{\gamma\gamma^{\prime}}\left(\theta\right)$ is $\left(T-1\right)K\times\left(T-1\right)K$, $H_{\alpha\gamma^{\prime}}\left(\theta\right)$ is $\left(N-1\right)K\times\left(T-1\right)K$, and $H_{\gamma\alpha^{\prime}}\left(\theta\right)=\left[H_{\alpha\gamma^{\prime}}\left(\theta\right)\right]^{\prime}$. The next lemma states that the quadratic form can be decomposed into two dominating terms of order $O(1/\sqrt{NT})$. They establish the two additive bias terms $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$, arising from estimating individual effects and time effects, respectively.
Lemma E.1
@@
$=-\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)+2\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right](1+o_{\mathbb{P}}(1)).$ (E.1)
@@
$\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right]=O\left(1/(NT)\right),$ (E.2)
@@
$\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right]=O(1/T),\qquad\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right]=O(1/N).$ (E.3)
@@
Under Assumptions 1 and 1, using Davydov’s inequality, we have $\sup_{i,j}\sup_{\theta\in\Theta}\operatorname*{Var}[h_{i,j}^{\alpha}(\theta)]=o(1)$ as $N,T\rightarrow\infty$, where $\operatorname*{Var}$ denotes the variance and $h_{i,j}^{\alpha}(\theta)$ is the $\left(i,j\right)$-th element of $N\cdot H_{\alpha\alpha^{\prime}}\left(\theta\right)$ with $i,j=1,\ldots,N-1$. Invoking the Markov’s inequality,
@@
uniformly for $\theta\in\Theta$, where $\|A\|_{\max}$ denotes the $\max$-norm of a matrix $A$. In addition, by matrix inversion in block form, it follows that
@@
For $i=1,\ldots,N-1$ and $t=1,\ldots,T-1$, define $h_{i,t}^{\ast\alpha\gamma}(\theta)$ as the $(i,t)$-th element of $H_{\alpha\gamma^{\prime}}^{\ast}(\theta)$. We first show (E.2). We have
@@
$=:A_{NT,1}(\theta)-A_{NT,2}(\theta)-A_{NT,3}(\theta)+A_{NT,4}(\theta).$ (E.4)
@@
$\quad\quad\times\sup\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\left[\left(1-\frac{m}{T-1}\right)[a_{i}(m)]^{1-2/\eta}\right]+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
@@
$\quad\quad\times\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\left[\left(1-\frac{m}{T-1}\right)[a_{i}(m)]^{1-2/\eta}\right]+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
@@
Therefore, (E.2) follows from (E.4).
@@
noting that $\sup_{i\neq j}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in $\theta$. Thus, the first part of (E.3) follows from (E.5). The second part of (E.3) follows from a similar procedure.
Combining (E.1), (E.2), and (E.3),
@@
$r^{s}\left(\theta,a\right)$ $:=\left(r_{\psi_{1}}^{s}\left(\theta,a_{1}\right),\ldots,r_{\psi_{(N+T-2)K}}^{s}(\theta,a_{(N+T-2)K})\right)^{\prime},$ (E.6)
@@
Proof. In the proof, we retain the notation and setup from Lemma E.1. From (2.6), it is evident that the dominant term in the expansion of $\widehat{\psi}(\theta)-\psi\left(\theta\right)$ is ${H}(\theta)]^{-1}s(\theta)$. Specifically,
@@
$=-\begin{pmatrix}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\\ H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\end{pmatrix}+[-\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\cdot o_{\mathbb{P}}(1)$ (E.7)
@@
$N\sqrt{T}u_{i}^{\alpha}(\theta)=N\sum_{j=1}^{N-1}\left[h_{i,j}^{\ast\alpha}(\theta)\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{j,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)\right].$ (E.8)
@@
$T\sqrt{N}u_{i}^{\alpha\gamma}(\theta)=\sum_{t=1}^{T-1}\left[h_{i,t}^{\ast\alpha\gamma}(\theta)\left(\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\left[\widetilde{s}_{j,t}^{\gamma}(\theta)-\widetilde{s}_{j,T}^{\gamma}(\theta)\right]\right)\right].$ (E.9)
@@
uniformly in $\theta$. From (E.8) and (E.9), we conclude that
$\sup_{\theta\in\Theta}\|H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(N\sqrt{T})\right),\qquad\sup_{\theta\in\Theta}\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(T\sqrt{N})\right).$ (E.10)
@@
$\sup_{\theta\in\Theta}\|H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(T\sqrt{N})\right),\qquad\sup_{\theta\in\Theta}\|H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(N\sqrt{T})\right).$ (E.11)
Thus, combining (E.7), (E.10 ), and (E.11), we have
@@
$\left\|\widehat{\psi}(\theta)-\psi(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\left(1/(NT)^{3/4}\right).$ (E.12)
From (E.12), we conclude that $\sup_{\theta\in\Theta}\|G_{\psi_{j}}\left(\theta,a_{j}\right)-\partial_{\psi_{j}}H(\theta)\|_{\max}=o_{\mathbb{P}}(1)$. Without loss of generality, let $\psi_{j}=\alpha_{i}$ for $i=1,\ldots,N-1$. We can express $\partial_{\alpha_{i}}H(\theta)$ as
@@
where $e_{i}$ is an $(N-1)\times 1$ vector with the $i$-th component equal to $1$ and all other elements equal to 0. Using (E.12), for each $j=1,\ldots,N-1$, we write
@@
$=:\frac{1}{2}\left\{C_{NT,1}(\theta)+C_{NT,2}(\theta)+C_{NT,3}(\theta)\right\}\left[1+o_{\mathbb{P}}(1)\right].$ (E.13)
Let $v_{i}^{\alpha}(\theta)$ denote the $i$-th element of $H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)$, for $i=1,\ldots,N-1$. Using (E.8) and (E.9), it can be shown that
@@
$=O_{\mathbb{P}}\left(1/(NT)\right).$ (E.14)
@@
$=O_{\mathbb{P}}\left(1/(NT)\right),$ (E.15)
@@
$r(\theta)$ $:=\frac{1}{2}(r^{s}(\theta,a))^{\prime}[H\left(\theta\right)]^{-1}r^{s}(\theta,a)+r^{l}(\theta,a_{0}),$ (E.16)
$r^{l}\left(\theta,a_{0}\right)$ $:=\frac{1}{3}\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}(\theta,a_{0}\times\iota_{(N+T-2)K})$ (E.17)
@@
Lemma E.3
Under Assumption 1, $r(\theta)$ defined by (E.16) satisfies
@@
Proof. In the proof, we retain the notation and setup from Lemmas E.1 and E.2. From Lemma E.2, we know that $NT\cdot r^{s}\left(\theta,a\right)$ is comparable to $(NT)^{3/4}\cdot s(\theta)$. Following similar arguments as in Lemma E.1, we can show that
@@
Moreover, from (E.12), (E.14), and (E.14), we have
@@
$=O_{\mathbb{P}}\left(1/(NT)^{5/4}\right).$ (E.18)
By Assumption 1 and applying the $L_{1}$ convergence theorem to E.18, we find that
@@
for $r\left(\theta\right)$ defined in (E.16). The claim in Theorem 1 follows directly from Lemmas E.1 and E.3.
@@
Proof of Theorem 2. We retain the notation introduced in Lemmas E.1-E.3. We first show
$\widehat{B}_{\alpha}(\theta)-B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),\qquad\widehat{B}_{\gamma}(\theta)-B_{\gamma}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$ (E.19)
To show the first part of (E.19), note that
@@
$\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)$ $:=\operatorname*{diag}\left\{\widehat{s}_{1}^{\alpha}\left(\theta\right)\left(\widehat{s}_{1}^{\alpha}\left(\theta\right)\right)^{\prime},\ldots,\widehat{s}_{N-1}^{\alpha}\left(\theta\right)\left(\widehat{s}_{N-1}^{\alpha}\left(\theta\right)\right)^{\prime}\right\},$
$\widehat{S}_{\alpha\alpha}^{2}\left(\theta\right)$ $:=(\iota_{N-1}\iota_{N-1}^{\prime})\otimes\left[\widehat{s}_{N}^{\alpha}\left(\theta\right)\left(\widehat{s}_{N}^{\alpha}\left(\theta\right)\right)^{\prime}\right].$
@@
$\quad\quad\times\operatorname*{diag}\left\{\widetilde{s}_{1,t}^{\alpha}(\theta)\left(\widetilde{s}_{1,s}^{\alpha}(\theta)\right)^{\prime}-\widehat{s}_{1,t}^{\alpha}(\theta)\left(\widehat{s}_{1,s}^{\alpha}(\theta)\right)^{\prime},\ldots,\widetilde{s}_{N-1,t}^{\alpha}(\theta)\left(\widetilde{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}-\widehat{s}_{N-1,t}^{\alpha}(\theta)\left(\widehat{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}\right\}$
@@
$\quad\quad\times\operatorname*{diag}\left\{\mathbb{E}\left[\widetilde{s}_{1,t}^{\alpha}(\theta)\left(\widetilde{s}_{1,s}^{\alpha}(\theta)\right)^{\prime}\right],\ldots,\mathbb{E}\left[\widetilde{s}_{N-1,t}^{\alpha}(\theta)\left(\widetilde{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}\right]\right\}.$ (E.20)
@@
given that $\tau/T\rightarrow 0$. Combining this with (E.20), we have
$\sup_{\theta\in\Theta}N^{2}\left\|S_{\alpha\alpha}^{1}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)\right\|_{\max}=o_{\mathbb{P}}(1/T).$ (E.21)
From the proof of Lemma E.2, we also have $\sup_{\theta\in\Theta}\|\widehat{\psi}(\theta)-\psi(\theta)\|_{\max}=o_{\mathbb{P}}(1)$. Moreover, under the conditional strong mixing assumption, each element of $T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ converges in probability to the corresponding element in $T^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ by the law of large numbers. Therefore (E.21) further leads to
@@
where the last equality follows from, as shown in Lemma E.2, $\sum_{j=1}^{N-1}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta$. The above leads to
$\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)\left[S_{\alpha\alpha}\left(\theta\right)-\widehat{S}_{\alpha\alpha}\left(\theta\right)\right]\right\}=o_{\mathbb{P}}(1/\sqrt{NT}).$ (E.22)
@@
$\operatorname*{trace}\left\{\left[\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)-H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\right]S_{\alpha\alpha}\left(\theta\right)\right\}=o_{\mathbb{P}}(1)\cdot\operatorname*{trace}\left\{H_{\alpha\alpha^{\prime}}^{\ast}(\theta)S_{\alpha\alpha}\left(\theta\right)\right\}=o_{\mathbb{P}}(1)\cdot B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$ (E.23)
Combining E.22 and E.23 leads to the first part of E.19. The second part of E.19 can be shown similarly.
@@
As from Lemma E.1, we have
@@
The conclusion follows by combining this and (E.19).
@@
$0=\triangledown_{\theta}l(\theta_{0})+\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0})+o_{\mathbb{P}}(1/\sqrt{NT})$ (E.24)
@@
Thus, using Assumption 2 and Slutsky’s theorem, we obtain from (E.24) that
$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})=-\left\{\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\right\}^{-1}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (E.25)
Given the conditional cross-sectional independence and the conditional strong mixing condition (Assumption 1), (E.25) leads to
$\left[\Sigma^{-1}\Omega\Sigma^{-1}\right]^{-1/2}\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}\left(0,\mathbb{I}_{K}\right)$ (E.26)
@@
Since we consider a likelihood model, $\Sigma^{-1}\Omega\Sigma^{-1}$ in (E.26) simplifies to $\Sigma^{-1}$ under the information matrix equality ($\Sigma=\Omega$).